OpenStudy (anonymous):

Where do theses two lines intersect? x^2+y^2=49 y=-2X+2 (I completed partial Work, don't fully understand matrix equation)

OpenStudy (anonymous):

Well, they aren't both lines. One is the equation of a circle with radius 7. Either way, solving this just requires you plug in y into the circle equation. This gives: \[x^2+(-2x+2)^2=49 \implies x^2+4x^2+4-8x=49 \implies 5x^2-8x-45=0\] Using the quadratic equation we find that: \[x= \frac{4 \pm \sqrt{241}}{5}\] Plug this into the y equation to solve for y.

OpenStudy (anonymous):